OpenStudy (anonymous):

Determine the number of solutions of 2x2 - 3x - 5 = 0. 0 1 (double root) 2

OpenStudy (goformit100):

1 (double root)

OpenStudy (anonymous):

thankss =)

OpenStudy (anonymous):

use the discriminant from the quadratic formula: b^2 - 4ac

OpenStudy (anonymous):

there are 2

OpenStudy (anonymous):

*2 solutions

OpenStudy (anonymous):

factor 2x^2 - 3x - 5 = 0 you get: (x+1)(2x-5)=0 solutions are x=-1 and x=5/2